# Sample-Based Conservative Bias Linear Power Flow Approximations

### v1

@@ IV Numerical results
In this section, we conduct numerical experiments on several test cases to examine the behavior of CBLA, highlight the benefits of error function design, and demonstrate the effectiveness of CBLA in a simplified OPF problem.
@@
The test cases used in the simulations are case6ww, case14, and the IEEE 24-bus system, all of which are accessible in Matpower [19]. For approximations of voltage and current flow, we draw $500$ samples by varying the power injections within a range of $70\%$ to $130\%$ of their nominal values. Both voltage and current flow values are reported in per unit (pu). We use the $\ell_{1}$ norm as the loss function $\mathcal{L}(\,\cdot\,)$. The numerical simulation was conducted in Julia using the Optim package.
IV-A Conservative Bias Linear Approximations
@@
We begin our numerical tests by examining the effects of changing the error function in (7) in our CBLA approach. As discussed in Section III-B, error functions are designed to balance conservativeness and accuracy. By testing different error functions, we aim to understand how they impact the number of violated samples and accuracy of the approximated power flow equations.
@@
In Fig. 4, we present the results of using the CBLA approach to intentionally overestimate the predicted current flow from bus 3 to bus 24 in IEEE 24-bus system using different quadratic error functions. The error functions employed in this test are defined as $g(\epsilon)=x^{2}$ and $h(\epsilon)=\alpha x^{2}$, where $\alpha$ is a parameter that we vary across the test. Specifically, we adjust $\alpha$ to take values of 1, 10, 100, and 1000. When $\alpha=1$, the error functions are equivalent (i.e., $g(\epsilon)=h(\epsilon)$), indicating that there is no difference in cost between violating and not violating conservativeness (i.e., the error function is the squared-$\ell_{2}$ loss). Under this condition, almost all samples are well approximated, but several samples fall below the zero approximation error line, indicating a deviation from the overestimating objective.
As we increase $\alpha$, fewer samples fall below the zero approximation error line, suggesting improved adherence to the overestimation goal. However, this improvement comes at the cost of lower overall accuracy. At the highest value of $\alpha=1000$, most samples are overestimated as intended, but the level of accuracy appears to be the lowest among all the plots. This trade-off highlights the importance of carefully selecting the value of $\alpha$ to achieve a suitable balance between overestimation and accuracy.
@@
TABLE I: Approximated current flow errors and number of violated samples at specific buses in IEEE 24-bus system
@@
The data presented in Table I illustrates the relationship between the number of violated samples and the average approximated current flow errors across different values of $\alpha$ at different lines. These results align with the trend observed in Fig. 5, confirming that as $\alpha$ increases, the average error per sample increases while the number of violated samples decreases.
IV-B Application: Simplified optimal power flow
While our main goal is to use our CBLA approach for challenge problems such as bilevel problems and mixed-integer nonlinear programs, these are outside the scope of this paper. Instead, we concentrate on showcasing results within a simplified optimal power flow (OPF) framework, which offers a conceptual demonstration and a foundation for comparing different linear approximations (DC power flow and conservative bias linear approximation (CBLA)). This simplified version of the OPF scenario imposes constraints on voltages at buses where $P$ and $Q$ are known, and power generation within defined ranges, excluding line flow limits (see [12] for the full problem setup).
TABLE II: Results from OPF comparing solutions from AC-, DC-, and CBLA-OPF
Formulation Case
case6ww case14
AC-OPF 2986.04 5368.30
DC-OPF 2995.15 (0.31%) 5368.52 (0.004%)
Violation $V$ (0.029 pu) No violation
CBLA-OPF ($\alpha=1$) 2987.28 (0.04%) 5368.52 (0.004%)
Violation No violation $V$ (0.004 pu)
CBLA-OPF ($\alpha=10^{2}$) 2987.42 (0.05%) 5368.52 (0.004%)
Violation No violation No violation
CBLA-OPF ($\alpha=10^{6}$) 2987.51 (0.05%) 5368.52 (0.004%)
Violation No violation No violation
Table II presents the outcomes of applying the AC power flow equations and various power flow approximations in a simplified optimal power flow (OPF) scenario. The AC-OPF serves as the baseline, with no violations and providing a reference for comparing the percentage difference in optimal cost against other formulations. In the overestimating CBLA approach, we utilize error functions $g(\epsilon)=\epsilon^{2}$ and $h(\epsilon)=\alpha\epsilon^{2}$, while for underestimating CBLA, we use error functions $g(\epsilon)=\alpha\epsilon^{2}$ and $h(\epsilon)=\epsilon^{2}$. As discussed in Section IV-A, setting $\alpha=1$ indicates that the error function does not differentiate between overestimating or underestimating the linearization error.
The results indicate that the DC-OPF leads to a maximum voltage violation of $0.029$ pu and an optimal cost that is $0.31\%$ higher than the AC-OPF’s solution for case6ww. In case14, the CBLA-OPF causes only a minor voltage violation of $0.004$ pu when $\alpha=1$. When $\alpha=10^{2}$ and $10^{6}$ (with the latter closely representing the conservative linear approximations in our previous work [12]), both test cases have no voltage violations. Moreover, the optimal costs using DC-OPF and CBLA-OPF are only $0.004\%$ higher than the optimal solution from AC-OPF in case14. These results suggest that the flexibility of CBLA allows us to select a suitable function that optimizes cost depending on the specific system being analyzed.
V Conclusion and future work
@@
This paper presents a conservative bias linear approximation (CBLA) approach for approximating the power flow equations. This approach strives to balance conservativeness and accuracy while maintaining linearity in the approximations. The numerical results highlight the potential advantages of using CBLA for power flow problems. By selecting an appropriate error function, we can achieve an effective balance between conservativeness and accuracy. Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions, ultimately enhancing performance and reliability in power system optimization.
@@
In our future work, we aim to extend our current approach by developing additional conservative bias approximations through the use of piecewise linearizations formulated as neural networks. Moreover, we plan to apply our proposed approach to a broader range of power system planning and resilience tasks. This includes tackling complex bilevel problems as well as conducting capacity expansion planning studies.
@@ Acknowledgement
The authors thank the Department of Mechatronics Engineering at Rajamangala University of Technology Phra Nakhon.

### v2

@@ V Numerical results
In this section, we conduct numerical experiments on several test cases to examine the behavior of CBLA, highlight the benefits of error function design, and demonstrate the effectiveness of CBLA in a UC problem.
@@
The test cases used in the simulations are the IEEE 24-bus system and the IEEE 30-bus system, all of which are accessible in Matpower [23]. For approximations of voltage and current flow, we draw $500$ samples by varying the power injections within a range of $70\%$ to $130\%$ of their nominal values. Both voltage and current flow values are reported in per unit (pu). We use the $\ell_{1}$ norm as the loss function $\mathcal{L}(\,\cdot\,)$. The numerical simulation was conducted in MATLAB using Yalmip [24] and in Julia using the Optim package.
V-A Conservative Bias Linear Approximations
@@
We begin our numerical tests by examining the effects of changing the error function in (7) in our CBLA approach. As discussed in Section III-B, error functions are designed to balance conservativeness and accuracy. By testing various error functions, we aim to understand their impact on the number of violated samples and the accuracy of the approximated power flow equations.
@@
In Fig. 5, we present the results of using the CBLA approach to intentionally overestimate the predicted current flow from bus 3 to bus 24 in IEEE 24-bus system using different quadratic error functions. The error functions used in this test are defined as $g(\epsilon)=\epsilon^{2}$ and $h(\epsilon)=\alpha\epsilon^{2}$, where $\alpha$ is a parameter that we vary across the test. Specifically, the parameter $\alpha$ modifies the bias in the CBLA formulation; a higher value of $\alpha$ increases the tendency toward conservativeness. We adjust $\alpha$ to take values of 1, 10, 100, and 1000. When $\alpha=1$, the error functions are equivalent (i.e., $g(\epsilon)=h(\epsilon)$), implying no distinction in cost between violating and not violating conservativeness (i.e., the squared-$\ell_{2}$ loss). In this scenario, most samples are well approximated, but some fall below the zero approximation error line, indicating a deviation from the overestimating objective.
Increasing $\alpha$ reduces the number of samples falling below the zero approximation error line, indicating better adherence to the overestimation goal. However, this improvement sacrifices overall accuracy. At $\alpha=1000$, where most samples are intentionally overestimated, the accuracy tends to be lower compared to other scenarios. This trade-off highlights the importance of carefully selecting the value of $\alpha$ to achieve a suitable balance between overestimation and accuracy.
@@
TABLE I: Approximated current flow errors and number of violated samples at representative buses in IEEE 24-bus system
@@
The data in Table I illustrates the relationship between the number of violated samples and the average approximated current flow errors across different $\alpha$ values at different lines. These results align with the trend observed in Fig. 6, confirming that as $\alpha$ increases, the average error per sample increases while the number of violated samples decreases.
V-B Application: Unit commitment
The comparison is conducted on the IEEE 30-bus test system, which has 6 generators. Voltage magnitude limits are set to 0.94 p.u. and 1.05 p.u. at all buses, and current magnitude limits are derived from the values in mpc.branch(:,RATE_A) in the MATPOWER case files. The cost of load shedding is $C^{\text{shed}}(L_{i,t})=500L_{i,t}$. We draw $1000$ samples by varying power injections, both loads and generations, to cover possible load shedding and generation variation within a range of $70\%$ to $130\%$ of their nominal values. As shown in Table II, we report both the total cost, which includes load shedding penalties, and the generation-only cost, which excludes them to better isolate the impact of load shedding. Although the DC-based formulation often yields lower generation-only costs, it leads to much higher total costs due to frequent and severe load shedding. As illustrated in Fig. 7, the generator commitment schedules from CBLA differ from those of the DC and first-order Taylor methods, highlighting CBLA’s ability to adapt commitment decisions based on a more accurate power flow approximation.
Figure 7: Generator commitment schedules for the 30-bus system using three different power flow linearization methods: DC, first-order Taylor, and CBLA (when $\alpha=10$ and $100$). Each row block corresponds to a generator across five scheduling hours. Blue indicates the generator is ON, while red indicates it is OFF.
This difference is especially evident in hour 4, where the DC-based approach leads to 11.82% load shedding and a total cost of \$20,417. In comparison, the first-order Taylor approximation reduces load shedding to 5.01%, with a much lower cost of \$11,683. CBLA reduces load shedding even further to just 1.92%, while lowering the total cost to \$8,053.7—clearly demonstrating the value of its adaptive linearization.
TABLE II: Hourly Comparison Across Different Approximations
H1 H2 H3 H4 H5
Load Demand (MW) 152.2 192.13 234.29 287.33 219.27
DC Load Shedding (%) 0% 4.37% 2.57% 11.82% 15.23%
Outputs (MW)
Gen 1 158.28 193.44 159.05 205.66 196.71
Gen 2 0 0 80 61.62 0
Gen 3 – Gen 6 0 0 0 0 0
Total Cost (\$) 1818 6328 6090 20417 18869
Gen-Only Cost (\$) 1817.6 2135.2 3076.1 3431.5 2167.3
First-Order Taylor Load Shedding (%) 0% 0% 0.70% 5.01% 2.42%
Outputs (MW)
Gen 1 52.84 85.27 134.71 202.56 189.72
Gen 2 68.48 80 80 62.73 0
Gen 3 0 0 0 0 0
Gen 4 33.86 32.41 26.96 22.16 33.90
Gen 5 – Gen 6 0 0 0 0 0
Total Cost (\$) 3483 3682 4800 11683 5875
Gen-Only Cost (\$) 3483.0 3682.1 3978.0 4480.5 3219.1
CBLA ($\alpha=10$) Load Shedding (%) 0% 0% 0% 1.92% 0%
Outputs (MW)
Gen 1 58.14 58.32 99.31 152.81 127.40
Gen 2 0 0 0 0 0
Gen 3 26.96 27.47 45.73 46.77 50.00
Gen 4 50.40 55.00 53.42 50.02 47.48
Gen 5 18.73 24.61 0 0 0
Gen 6 0 29.04 40 40 0
Total Cost (\$) 4506.2 5660.4 4929.7 8053.7 3958.8
Gen-Only Cost (\$) 4506.2 5660.4 4929.7 5299.5 3958.8
CBLA ($\alpha=100$) Load Shedding (%) 0% 0% 0% 1.92% 0.93%
Outputs (MW)
Gen 1 110.53 110.45 99.31 152.81 100.66
Gen 2 0 0 0 0 80.00
Gen 3 45.01 45.40 45.73 46.77 42.51
Gen 4 0 0 53.42 50.02 0
Gen 5 0 0 0 0 0
Gen 6 0 40.00 40.00 40.00 0
Total Cost (\$) 2637.0 3799.1 4929.7 8053.7 4830.8
Gen-Only Cost (\$) 2637.0 3799.1 4929.7 5299.5 3811.4
While the DC and first-order Taylor methods rely on a single linearization of nonlinear power flows, CBLA builds a sample-based conservative approximation over a defined operating region, delivering significant improvements. Across all five hours, CBLA eliminates load shedding in four hours and limits it to minimal levels in the most challenging hour. CBLA consistently achieves the lowest total cost. For example, in hour 5, CBLA attains a lower total cost of \$3,958.8 compared to \$5,875 with the Taylor approximation and \$18,869 with the DC approximation.
The table also includes results for a more conservative CBLA variant with $\alpha=100$. With $\alpha=100$, the UC solution tends to avoid constraint violations. However, this added conservativeness comes at the cost of reduced accuracy in approximations. As a result, when the generation decisions from the UC problem are passed to the AC OPF, the resulting solution may still exhibit load shedding—particularly in Hour 5, where 0.93% of the demand remains unmet.
This trade-off underscores the need to balance conservativeness and approximation accuracy. CBLA’s tunable parameter $\alpha$ allows us to adjust this balance to achieve the desired feasibility and flexibility. By tailoring the approximation to system conditions, CBLA helps reduce constraint violations and yields UC solutions that generally remain feasible under the nonlinear AC power flow model.
Moreover, the benefits of CBLA become more pronounced in larger systems or under tight feasibility margins. Unlike the first-order Taylor approximation, which is inherently local, CBLA is constructed to remain conservative across a sampled region of interest. This range-aware property improves cost-efficiency in decision-making. Compared to traditional methods, CBLA provides a more adaptive solution framework that maintains the tractability of linear programming while reducing the gap between the models of linear and nonlinear power systems.
VI Conclusion and future work
@@
This paper presents a conservative bias linear approximation (CBLA) approach for approximating the power flow equations, aiming to balance conservativeness and accuracy while preserving linearity. The numerical results highlight the potential advantages of using CBLA for power flow problems. Selecting suitable error functions enables an effective balance between conservativeness and accuracy. Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions. When applied to the unit commitment problem, the CBLA-based formulation consistently outperformed formulations based on DC power flow and first-order Taylor approximations, achieving lower overall operating costs in the subsequent evaluation of the AC optimal power flow.
@@
In our future work, we aim to extend our current approach by developing additional conservative bias approximations through the use of piecewise linearizations. Moreover, we plan to apply our proposed approach to a broader range of power system planning and resilience tasks. This includes tackling complex bilevel problems, conducting capacity expansion planning studies, and extending the application of CBLA to unbalanced three-phase systems.
@@ Acknowledgement
P. Buason thanks the Department of Mechatronics Engineering at Rajamangala University of Technology Phra Nakhon.
Appendix: First-Order Taylor Approximation of Voltage and Current Constraints
Suppose the power injections $\bm{P}$ and $\bm{Q}$ deviate from their nominal values $\bm{P}^{\text{nom}}$ and $\bm{Q}^{\text{nom}}$. Let $\mathcal{B}^{\text{PQ}}$ denote the set of PQ buses, $\widetilde{V}_{i,t}$ denote the first-order Taylor approximation of the voltage magnitude at bus $i$ and time $t$, and let the abbreviation “nom” indicate that a quantity is evaluated at the nominal operating point, which is computed separately for each hour based on the corresponding loads. For each $i\in\mathcal{B}^{\text{PQ}}$, the first-order Taylor approximation of the voltage magnitude is:
$\widetilde{V}_{i,t}=V_{i}^{\text{nom}}+\frac{\partial V_{i}}{\partial\bm{P}}\bigg{|}_{\text{nom}}\cdot(\bm{P}_{t}-\bm{P}^{\text{nom}})+\frac{\partial V_{i}}{\partial\bm{Q}}\bigg{|}_{\text{nom}}\cdot(\bm{Q}_{t}-\bm{Q}^{\text{nom}}),$ (15)
where the partial derivatives $\frac{\partial V_{i}}{\partial\bm{P}}$ and $\frac{\partial V_{i}}{\partial\bm{Q}}$ are evaluated at the nominal operating point using the inverse of the power flow Jacobian. For non-PQ buses, the voltage magnitude is fixed at its nominal value:
$\widetilde{V}_{i,t}=V_{i}^{\text{nom}},\quad\forall i\notin\mathcal{B}^{\text{PQ}}.$ (16)
To model current flow limits, we approximate the real part of the power flow on each transmission line $(i,k)\in\mathcal{L}$ using the difference in bus voltage magnitudes from (16):
$\Re(S_{ik,t})\approx|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t}),$ (17)
where $\Re(\;\cdot\;)$ denotes the real part of a complex quantity and $|Y_{ik}|$ is the magnitude of the admittance between buses $i$ and $k$. We then enforce the following absolute value inequality to approximate current magnitude limits:
$-I_{ik}^{\text{max}}\leq|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t})\leq I_{ik}^{\text{max}},\;\forall(i,k)\in\mathcal{L},\;\forall t\in\mathcal{T}.$ (18)
To allow controlled violations of these approximations similar to (11)–(12), slack variables may be added as needed in the optimization formulation.